A 90-kg football player stands on a frictionless lake of ice. He catches a 3-kg football and then moves at 400 cm/s. What was the initial velocity of the football? 


1
Expert's answer
2021-04-02T10:33:53-0400

"mv_0=(m+M)v\to v=\frac{(m+M)v}{m}=\frac{(3+90)\cdot4}{3}=124\ (m/s)" . Answer